OpenStudy (anonymous):

solve .. 4b/9 - b/5 = 11

OpenStudy (anonymous):

\[\frac{4b}{9}-\frac{b}{5}=11\] Get the LCM of 9 and 5 and mutliply the numerator with 9 and 5 \[\frac{5(4b)-9(b)}{45} = 11\] \[5(4b)-9(b) = 45\times11\]

OpenStudy (anonymous):

hope u can do the rest!

OpenStudy (anonymous):

nah , im lost .. lol

OpenStudy (anonymous):

okay... 5(4b)−9(b)=45×11 u understnd this prt?

OpenStudy (anonymous):

u following?